OpenStudy (anonymous):

Write a rule for the linear function with the following description: the graph has slope 2 and contains the point (-1, 1): A. y = 2x - 1 B. y = 2x + 3 C. y = x + 1 D. y = -x + 2

OpenStudy (texaschic101):

y = mx + b m is your slope...so m = 2 (-1,1)...x = -1 and y = 1 now we sub and solve for the y int (b) 1 = 2(-1) + b 1 = -2 + b 1 + 2 = b 3 = b so your equation is : y = 2x + 3 questions ?

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

sure thing :)